OpenStudy (anonymous):

How would i find the square root of 3.344x10^-27? How about 8.36x10^-28? By that I mean if i had the following problem: 1.881x10^-11=8.36x10^-28 x v1f^2 + 3.344x10^-27 x v2f^2 and v1f=150000000 - 4v2f I would get 1.881x10^-11=8.36x10^-28 (15x10^7 - 4v2f) + 3.344x10^-27 x v2f^2 Therefore how would I write down the square root of both sides?

OpenStudy (anonymous):

By the way v1f and v2f are the final velocities. It's a physics question but i've forgotten how to do it somehow

OpenStudy (anonymous):

\[1.881\times 10^{-11}=8.36\times 10^{-28} \times v1f^2 + 3.344\times 10^{-27} \times v2f^2\] is the is the problem? what is the variable?

OpenStudy (anonymous):

\[1.881\times 10^{-11}=8.36\times 10^{-28} \times ( 150000000 - 4v2f)^2+ 3.344\times 10^{-27} \times v2f^2\]

OpenStudy (anonymous):

Nevermind I got it now. All i wanted to know was how to find something like \[\sqrt(8.36×10−28×v1f2\] and because i was tired i forgot how to do it. thanks for replying